OpenStudy (anonymous):

The formula to compute a person's body mss index is B=703 x w/h squared. Solve the formula for w.

OpenStudy (anonymous):

i need this today.

OpenStudy (anonymous):

BMI = 703w/h² You're given the inputs , so to find the BMI, it's just plug-and-solve: BMI = 703(200)/(72²) BMI = 140600 /.5184 BMI = 27.1 (rounded to tenths) If you want to find the weight required to get a specific BMI, then plug in the known values and solve for what y ou're looking for. In this case: w 26 = 703w/(72²) 26 = 703w/5184 134784 = 703w w = 191.7 So he's have to lose 8.3 lbs (200 - 191.7)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

your welcome